### v1

# Early Jamming Detection in Mobile Indoor Scenarios via Deep Learning

Abstract
The current state of the art on jamming detection relies on link-layer metrics. A few examples are the bit-error rate, the packet delivery ratio, the throughput, and the increase of the signal-to-noise ratio. As a result, these techniques can only detect jamming ex-post, i.e., once the attack has already taken down the communication link. These solutions are unfit in mobile scenarios, e.g., drones, which might lose the link to the remote controller, being unable to predict the attack.
Our solution is rooted in the idea that a drone flying against a jammed area is experiencing an increasing effect of the jamming. Therefore, drones might use this phenomenon to detect jamming early, i.e., before it completely disrupts the communication link. Such an approach would allow drones and possibly their pilots to make an informed decision and maintain full control of the navigation, providing security and safety.
In this paper, we propose Bloodhound$+$, a solution for early jamming detection on mobile devices. Our approach analyzes raw physical-layer information (I-Q samples) acquired from the channel. We assemble this information into grayscale images, and we use sparse autoencoders to detect image anomalies caused by jamming attacks. To test our solution against a wide set of configurations, we acquired a large dataset of indoor measurements using multiple hardware, jamming strategies, and communication parameters. Our results indicate that Bloodhound$+$ can detect indoor jamming up to 20 meters away from the jamming source at the minimum available relative jamming power, with a minimum accuracy of 99.7%. Our solution is also robust to various sampling rates adopted by the jammer, as well as to the type of signal used for jamming.
Index Terms:
Wireless Security; Artificial Intelligence for Security; Drones Security.
I Introduction
The adoption of remotely piloted and (semi) autonomous robots and drones in everyday business is increasing at a high pace. Today, such technologies are increasingly used both outdoors, e.g., for disaster management, search and rescue, and goods delivery [1], and indoors, e.g., for inventory management, intra-logistics of items, inspection and surveillance [2]. In this context, leading market analysis companies estimate a current market value of up to 100.37 billion USD by 2029, with a compound annual growth rate of 25.5% in the next years [3].
As their role becomes more central, drones are increasingly the main target of several cybersecurity attacks. In particular, due to the reliance on wireless channels for communication, video streaming, and telemetry, the easiest way attackers can disrupt drones’ operation is through jamming attacks [4]. Jamming can significantly disrupt wireless communications in a given area through injection of high-power noise into the same channel used by legitimate communication parties [5]. In such cases, depending on the firmware onboard, the drone might return to the starting point of the mission, land, or even crash, with potential hazards for people in the area, especially for indoor applications [6].
At the time of writing, the most available solutions for jamming detection work by identifying sudden drops in the Bit-Error Rate (BER), the Packet Delivery Rate (PDR), or the Signal to Noise Ratio (SNR) of the communication link. Some of them also use the spectrogram of the signal in the Physical (PHY) layer, looking for sudden anomalies (see Section VI for an overview). Such solutions work reliably and effectively when applied to static deployments. However, they can detect jamming mainly ex-post, i.e., once jamming has already disrupted the regular operations of the communication link. The application of such approaches on drones would trigger the default actions listed above, possibly causing hazards to the drone and the surrounding people. In this context, mobile devices such as drones typically experience an increasing effect of the jamming (high BER, low PDR, low SNR) while approaching the jammed area. Drones might exploit such phenomena to deploy a solution able to detect jamming earlier, i.e., before entering an area where the high-power noise injected by the jammer completely disrupts the communication link. Such a solution is especially critical for indoor applications, where people and drones often work together in limited space.
Contribution. In this paper, we propose BloodHound$+$, an innovative solution for early jamming detection in mobile indoor scenarios leveraging state-of-the-art Deep Learning (DL) techniques. BloodHound$+$ allows to carry out jamming detection by converting raw PHY data, that is, I-Q samples, into images while detecting anomalies in their shape by resorting to autoencoders.
When applied to mobile entities, for example, autonomous or remotely piloted vehicles and drones, BloodHound$+$ can detect the approach of a jammed area well before jamming could completely disrupt the communication link, allowing the remote entity to maintain full control. To verify the effectiveness of our proposed approach, we carried out an extensive measurement campaign, involving multiple hardware devices (multiple Ettus Research X310 and LimeSDR radios), communication link configurations, and jamming conditions. Using such measurements, we tested the effectiveness of BloodHound$+$ and other competing approaches for the early detection of jamming. Our results show that BloodHound$+$ can detect jamming well earlier compared to benchmark solutions, for example, with an accuracy of $0.997$ when the adversary jams at a distance of $10$ m from the target with a Relative Jamming Power (RJP) of $0.1$. Our solution is also very robust to: (i) the distance from the jammer, (ii) the training set size, (iii) the number of acquired samples, (iv) the sampling ratios at the jammer and the receiver, (v) the type of jamming signal (tone, Gaussian, or deceptive), as well as, (vi) the adoption of different jamming hardware and radio types.
We note that this contribution extends and completes our previous work published in [7], by providing the following new content.
  • We focus on an indoor scenario, providing a brand new range of data considering additional hardware, modulation techniques, and jamming strategies.
  • We consider a stronger adversary model, assuming that the adversary is aware of the sampling rate and the modulation techniques used by the legitimate communication link. Such knowledge allows the adversary to optimize the parameters of the jamming attack, in order to boost its effectiveness and avoid detection at the same time.
  • We design a new optimized methodology for jamming detection, based on a one-class classifier of black and white images extracted from raw I-Q samples using sparse autoencoders.
  • We experimentally compare our new methodology with the ones proposed in [7] and [8], showing remarkable performances and improvements with reference to the RJP at the receiver, distance from the jammer, number of I-Q samples per image, training set size, and invariance to the hardware used for training.
  • We provide additional results on a new dataset, gathered using new hardware, namely, the LimeSDR.
  • We provide new results obtained on the data gathered by using various sampling rates at the jammer and receiver.
  • We provide new results on the usage of a new jamming strategy, i.e., deceptive jamming using the same modulation ( Binary Phase-Shift Keying (BPSK)) and signal of the legitimate communication link.
Roadmap. The rest of this paper is organized as follows. Section II introduces preliminary notions; Section III describes the scenario and adversarial model; Section IV provides the rationale and details of BloodHound$+$; Section V discusses our extensive measurement campaign and performance assessment of our solution, and finally Section VII draws the conclusion and outlines future work.
II Preliminaries
In this section, we introduce preliminary notions that are useful to the readers of this manuscript, i.e., digital modulation techniques (Section II-A) and autoencoders (Section II-B).
II-A Digital Modulation
Digital modulation schemes adopted in wireless communication systems preprocess baseband signals to make them suitable for transmission at high frequencies [9]. Typically, modulation techniques divide the bitstream to be transmitted into two orthogonal components, namely the I vector and the Q vector, linked in a complex value of type $I+jQ$, where the I vector is the real component and the Q vector is the imaginary component. Due to their orthogonality, such components can be transmitted together on the wireless channel without interfering and can also be recovered and assembled at the receiver to reconstruct the original bit-stream. In this context, a usual way to represent complex I-Q signals is through the I-Q plane, as shown in Fig. 1.
Figure 1: I-Q plane of a BPSK modulation. The receiver expects two symbols, i.e., $\left[i=-1,q=0\right]$ and $\left[i=1,q=0\right]$ (red crosses). However, due to the wireless channel, the received symbols are displaced (light grey area).
In particular, the number of expected I-Q values at the receiver (denoted as $n$) indicates the amount of bits that can be recovered through a single complex I-Q value. In general, we can recover $\log_{2}{n}$ bits from $n$ symbols, and thus, with reference to Fig. 1 showing the I-Q plane of a BPSK modulation, we can recover $n=2$ bits, i.e., $\left[i=-1,q=0\right]$ (b=0) and $\left[i=1,q=0\right]$ (b=1). We also notice that the values of the I-Q samples at the receiver always differ from those at the transmitter because of noise introduced by the hardware components of the devices and the fluctuations of the wireless channel between the transmitter and the receiver. To recover the original transmitted symbol, the receiver associates to the received I-Q sample the expected symbol whose distance from the received one is the shortest. Thus, the higher the impact of noise, the higher the chance that a received I-Q sample is associated with the wrong expected symbol, leading to an error and therefore to a higher BER. At the same time, with a given noise profile affecting the communication channel, the higher the modulation order $n$, the higher the amount of expected symbols and, thus, the higher the BER. This is the rationale for the use of lower-order modulation schemes for communication links affected by high levels of noise, such as satellite transmissions and mobile indoor applications. The intuition driving our work is that the collective displacement of I-Q samples from the expected one can be used to discriminate the presence of various levels of intentional interference, i.e., jamming, affecting the communication link. We will provide more details about our approach in Section IV.
II-B Autoencoders
Without loss of generality, autoencoders are a special type of Artificial (ANN) which can be trained to reconstruct their input [10]. Formally, the problem autoencoders solve is to find an encoder $A:\mathbb{R}^{d}\to\mathbb{R}^{p}$ and decoder $B:\mathbb{R}^{p}\to\mathbb{R}^{d}$ satisfying Eq. 1.
$\arg\min_{A,B}E[\Delta(\mathbf{x},B\circ A(\mathbf{x}))],$ (1)
where the symbol $\circ$” represents the composition operator, i.e., $B\circ A(\mathbf{x})=B(A(\mathbf{x}))$, $E$ represents the expectation of the distribution of the input $\mathbf{x}$, ’$A(\mathbf{x})$’ the encoded version of the input, known as the bottleneck of the autoencoder when $p<d$, and finally $\Delta$ is the reconstruction loss function, which measures the distance between the input of the ANN and the reconstruction of the input [10]. For our purposes, in line with many scientific contributions such as [10] and [11], $\Delta$ is the mean-squared-error (mse) function, as defined in Eq. 2 on two reference distributions $x$ and $y$.
$\text{mse}(\mathbf{x},\mathbf{y})=\frac{1}{d}\cdot\|\mathbf{x}-\mathbf{y}\|^{2}_{2}\text{ for all }\mathbf{x},\mathbf{y}\in\mathbb{R}^{d},$ (2)
being $d=M\cdot N$.
Traditionally, autoencoders have been used mostly for image generation and, in particular, for creating sets of images similar to the input ones. However, in the cybersecurity research domain, they are mainly used for anomaly detection. Let $c$ be an autoencoder trained on samples taken from a probability distribution $P$. Next, let $Q$ be a probability distribution such that $P\neq Q$. Then, we expect $c$ to have a smaller reconstruction error when tested on unseen samples from $P$ than when tested on unseen samples from $Q$. Therefore, the magnitude of the reconstruction error of $c$ on an unseen sample is a measure of the probability that such an unseen sample is not sampled from $P$, but from another distribution [11]. Consequently, we can define a specific error value $\tau$ as a classification boundary. All samples in which the auto-encoder $c$ makes a higher error than $\tau$ can be classified as being ’not from $P$’. In the literature, $\tau$ is often referred to as a threshold [12, 13].
In this work, we use autoencoders to build a statistical profile of the channel experienced between the mobile transmitter and the receiver under regular operating conditions. We provide more details in Section IV.
III System and Adversary Model
Fig. 2 shows the scenario and adversary model considered in this work.
Figure 2: Reference System and Adversary Model. While operating indoors, a drone tries to detect jamming (red color) early, before it might affect the quality of the legitimate communication link (blue color).
We consider an indoor scenario, where drones operate to achieve (semi)-automatized tasks, e.g., warehouse inspection. We consider the existence of a communication link between the drone and a Ground Control Station (GCS), which can be used either to pilot the drone, in the case of a Remotely-Piloted Aircraft System (RPAS), or to report telemetry data, in the case of (semi) autonomous operations. Independently of the usage of the communication link, we do not make any assumption on the nature of the drone operations, which can be either instructed by a human pilot or (semi) autonomous. Without loss of generality, we consider that the communication link between the drone and the GCS adopts the BPSK modulation scheme. This is a reasonable assumption, as such a scheme allows one to mitigate as much as possible the noise affecting indoor communication channels, being used also in modern WiFi standards.
We also consider the deployment of a static jammer in the area, which injects noise into the wireless channel used for communication between the drone and the GCS. We assume that such a jammer continuously emits interfering signals, with the highest possible transmission power, to affect as much as possible ongoing wireless communications in the deployment area. Unlike the contribution in [7], we do not make any assumptions about the specific jamming signal: it can be Additive White Gaussian Noise (AWGN), a single tone, and even deceptive jamming, adopting the same modulation scheme used by the legitimate communication link. Also, being possibly unaware of the sampling rate of the legitimate communication link, the jammer transmits signals with the highest possible sampling ratio, limited only by the hardware used to carry out the attack. Note that when the attacker is unaware of the modulation used by the legitimate communication link, he/she can perform modulation-agnostic jamming, e.g. by injecting AWGN or a single tone, centered on the channel of interest. Instead, if the attacker is aware of the modulation used by the target link, it can use deceptive jamming, i.e., injecting a signal characterized by the same digital modulation (and possibly the same message pattern) of legitimate messages, further complicating the detection process.
Being bounded by the maximum achievable transmission power, the jammer significantly impacts ongoing communications only in a specific area around its location. In fact, the Received Signal Strength (RSS) associated with the jamming signal depends on the distance between the jammer and the receiver location (drone), which is highest in the proximity of the jammer and decreases with further movement. The described wireless propagation effect generates a jammed area, where wireless communication is disrupted. In this context, the mobile receiver (drone), moving toward the jammed area, wants to promptly detect the jamming signal, before the effect of the jamming on the quality of the communication link becomes noticeable, i.e., it causes a significant increase in BER. In fact, such an early jamming detection mechanism would improve drone situational awareness, as it would allow the GCS to be aware of the imminent jamming and take action immediately, without relying on a predefined set of actions (e.g., landing, returning to the starting point).
IV Methodology
This section describes BloodHound$+$, i.e., the methodology we propose to detect jamming before it significantly affects the BER of the legitimate communication link significantly. In summary, BloodHound$+$transforms jamming detection into an anomaly detection problem on images generated by encoding the current state of the legitimate communication channel. Overall, we can identify two main building blocks of our solution: the Image Generation and the Jamming Detection, described below. Tab. I summarizes the main notation used below, with a short description.
TABLE I: Notation and brief description.
Notation Description
n
Number of samples per image.
M, N
Dimensions of the input image.
$a_{m,n}$
Generic pixel of the input image.
$d$
Dimension of vectors within autoencoders, with $d=M\cdot N$.
K
Encoder units.
J
Decoder units.
$\tau$
Autoencoder threshold value.
$MSE_{train}$
mse value obtained at training time.
Image Generation. Fig. 3 provides a graphical overview of the image generation process used in BloodHound$+$.
Figure 3: Graphical overview of the image generation process of BloodHound$+$. We acquire raw I-Q data through a generic SDR, we plot them through the traditional I-Q plane representation, we compute a bi-variate histogram based on the density of samples in given areas of the plane, and then, we obtain an image.
The input to the process is represented by a sequence of raw I-Q samples. Such samples can be collected using a SDR or any hardware capable of obtaining PHY wireless channel information (e.g., spectrum analyzers). The amount of samples used to generate images, namely $n$, is one of the degrees of freedom of our solution, and it can be configured by trading off the overall accuracy with the computational requirements of the solution (see Section V for a detailed evaluation of the impact of this parameter). We represent the sequence of I-Q samples through the traditional I-Q plane, where we display the component $I$ on the x-axis and the component $Q$ on the y-axis. In a benign scenario (no jamming), such a representation generates a number of clouds of I-Q values, approximately centered on the value of the expected symbol, as explained in Section II-A. Based on such a representation, we build a bivariate histogram. Specifically, we divide the I-Q plane around the cloud of points into tiles $N\times M$, where the values of $N$ and $M$ depend on the dimensions of the images we want to obtain. Then, for each tile $a_{m,n}$, we evaluate the number of I-Q samples that fall into the tile itself. When the number of I-Q samples falling on a tile exceeds the value $255$, we truncate it to the maximum value, to guarantee $a_{m,n}\in[0,255],\forall(m,n)$. We consider the output of such a process as a pixel value. As a consequence, the output of the image generation process is a grayscale image corresponding to the received profile of I-Q samples. Note that, in principle, we can also choose to work with color images (3-D matrices). In Section V we evaluate this configuration, adopted in [8], and show its pros and cons in the context of physical-layer jamming detection.
Jamming Detection. The jamming detection process is the building block of BloodHound$+$ dedicated to the timely detection of possible jamming affecting the wireless communication channel. It is a DL-based process using sparse autoencoders, so involving a training and testing process.
Fig. 4 shows the architecture of the adopted autoencoder.
Figure 4: Autoencoder architecture. We use a logsig encoder transfer function and a purelin decoder transfer function, with a total number of two hidden layers, a sparsity regularization term of 0.05, and the mse as a loss function, coming up with a sparse nonlinear autoencoder.
The input to the autoencoder is represented by the images obtained as a result of applying the image generation process to I-Q samples collected from the wireless channel. At training time, we acquire I-Q samples corresponding to the typical behavior of the communication link. We denote images corresponding to such a scenario configuration as unjammed images. We acquire $n$ I-Q samples and generate from them images of size $M\times N$. In our deployment, $n=10^{5}$ and $M=N=224$, to match the size of images used as input to various NN (we compare our performance to two benchmark solutions using Convolutional (CNN) in Section V). As encoder, we used the logarithmic sigmoid functions, with $K=16$ units (a.k.a. neurons). As a result, we obtain a latent representation of the input image, consisting of $K=16$ dimensions. Such a latent representation summarizes the relevant features of the input image, significantly compressing its dimension (compared to the input image). Then we submit the latent representation vectors to the decoder, using a linear decoder transfer function with a total number of $J=50,176$ neurons. In our context, using the logistic sigmoid activation function in the decoder’s units did not allow our solution to converge to a good solution. We suspect vanishing gradients to be the cause, which is further supported by the better convergence we obtain when using linear activation functions in our decoder layer. In principle, autoencoders allow the use of multiple hidden layers. Overall, the higher the number of hidden layers, the better the performance of the classifier at run-time, but also the higher the computational overhead of the methodology. Here, we use two hidden layers and the sparsity regularization technique, in line with the architecture of sparse nonlinear autoencoders. As we show in Section V, such a choice allows us to obtain remarkable classification accuracy, while achieving a computational cost lower than that of more complex architectures.
This process provides us with a reconstructed image of the same dimension as the input. We first convert the matrix of dimension $m\times n$ in a vector of dimension $d=M\cdot N$, concatenating the rows of the image one after the other. Then, we compute the mse loss function as in Eq. 2 (see Section II-B). During training, we acquire several images corresponding to the regular (expected) behavior of the wireless channel, building a corresponding profile of such a channel when displayed through images. We compute a threshold $\tau$ on such a profile, as explained in Section II-B, to distinguish the regular channel conditions from the unexpected one. At testing time, we compare the mse value obtained from a run-time acquisition of the wireless channel with the threshold $\tau$ previously cited. If the mse of the input image is equal or greater than the threshold, the autoencoder produces a positive prediction, meaning that there is a jammer that affects the communication channel. Otherwise, when mse is lower than the threshold, the autoencoder outputs a negative prediction, meaning there is no jamming. We provide below some more details on the training of the autoencoder and the threshold selection process of BloodHound$+$.
Training the autoencoder. For training the autoencoder, we use only unjammed images, i.e., images generated from I-Q samples acquired when there is no jamming affecting the communication link. We do not use any jammed images, i.e., images obtained from a jammed communication channel. We applied this strategy mainly since jamming can be performed in many different ways, typically unknown and unpredictable to the legitimate parties. Instead, our intuition is that we can build a profile of the expected conditions of the communication link, even in very noisy scenarios, and detect jamming as a deviation from such expected conditions. In this context, to guarantee reliable operations for the autoencoder, it is important to gather I-Q samples that cover the largest possible set of expected conditions of the communication channel. In fact, the reliability of the auto-encoder in identifying anomalies leading to jamming depends on the variety of conditions affecting the communication channel, thus reducing false positive events.
Threshold selection. Optimal selection of the decision threshold of an autoencoder is an actively researched problem, which does not have a universally optimal solution [14, 13]. In this work, we adopt the approach suggested by the authors in [12], i.e., we compute the threshold according to Eq. 3.
$\tau={mean}({MSE}_{\text{train}})+3.5\cdot{std}({MSE}_{{train}}),$ (3)
being ${MSE}_{{train}}$ the set of mses that the autoencoder compute on the training data, ${mean}$ the statistical average and ${std}$ the standard deviation. As demonstrated by the authors in [12] and confirmed by the authors in [13], such a choice is reasonable in scenarios where no anomalous samples are used in the training phase, as in our scenario. Moreover, as acknowledged by the authors in [13], this choice allows us to reduce false negatives when compared to the common choice, i.e., setting $\tau$ to the maximum mse observed on the training samples. In turn, such a choice allows us to boost the chances of detecting jamming earlier. Figure 5 reports an example of the threshold selection process on real data acquired with jammer jamming with $RJP=0.6$ at a distance of $10$ meters from the receiver (see Section V-A for details). We notice that the distribution of the mse values for unjammed images is characterized by smaller values compared to the one of jammed images, with only minimal overlapping on the tails of the distributions. Setting the threshold according to Eq. 3 allows one to reduce false negatives, without affecting performance.
Figure 5: Sample distribution of the mse of unjammed and jammed images and optimal threshold selection. The reported data refer to a jammer jamming with $RJP=0.6$ at a distance of $10$ meters from the receiver.
Finally, note that deploying an autoencoder for image-based jamming detection involves setting optimal values of the hyperparameters of such a tool. We discuss the selection of auto-encoder hyperparameters in Section V.
V Experimental Assessment
In this section, we provide the details of our extensive experimental assessment, carried out to evaluate BloodHound$+$ in real indoor scenarios. We introduce the actual measurements used for the following analysis in Section V-A, while in Section V-B we describe the experimental settings. Then Section V-C reports the performance of our approach and compares it with two solutions available in the literature. We extend such results further, evaluating BloodHound$+$ with different hardware and various channel sampling rates (Section V-D), and investigating the capability of identifying deceptive jamming (Section V-E).
V-A Measurements
In this paper, we build on top of the data provided as part of our contribution in [15] and extend such a dataset with new measurements obtained with new hardware and different configuration parameters.
All of the measurements discussed below have been acquired in an indoor office environment during a regular working day, with people moving around and possibly across the legitimate communication link. Such conditions match, as much as possible, those described in Section III.
The general setup of our measurements includes three entities, i.e., a transmitter, a receiver, and a jammer. In the first set of experiments, we considered SDR Ettus Research X310 [16] featuring a daughterboard UBX160 as reference hardware for all the entities involved in our measurement campaign. Here, we placed the transmitter and the jammer close to each other, while moving the receiver at various distances from the transmitter (see below). For the second set of experiments, we used different hardware, that is the LimeSDR [17]. It is a low-cost, open-source, SDR platform that can be used to support any wireless communication standard. For these experiments, we placed the transmitter and receiver $3$ meters away from each other and placed the jammer between them, at a distance of $1.5$ meters from both entities. We connected the SDRs either via Ethernet (Ettus X310) or USB 2.0 (LimeSDR) to two laptops, one controlling the data transmission and jamming processes and the other taking care of data reception. The received I-Q samples were stored on the laptop connected to the receiving SDR, and subsequently, uploaded to a centralized server for data analysis. Specifically, we used the High Processing Cluster (HPC) available at TU/e, Eindhoven, The Netherlands, providing a CPU E2124 with 4 cores running at 3.3 GHz and 32 GB of RAM, as well as 2 GPUs Tesla V1000 running at 32 with 256 GB of RAM.
Regarding the software, for both setups, we used the GNURadio v3.8 development toolkit [18]. We set the carrier frequency $f_{c}=900$ MHz for both the legitimate communication link and the jamming. The transmitter and receiver are configured to exchange packets containing a repeating sequence of 256 bytes, encoded by a Constellation Modulation block using the regular BPSK modulation scheme.
For the first setup using the SDR Ettus X310, we configured a sample rate of 1M samples per second at the transmitter, the receiver, and the jammer. We set the normalized transmission power and receiver gain to the maximum value of $1$, corresponding to approximately 15 dBm (32 mW) of transmission power. At the receiver, we set up the reception chain of the BPSK modulation, including (i) an Adaptive Gain Control (AGC) block, to mitigate the signal level fluctuations introduced by the multipath fading; (ii) a Symbol Sync, which performs timing synchronization; (iii) a Costas Loop, which locks to the center frequency of the signal and down-converts it to baseband; and, (iv) finally, a Constellation Decoder block, which decodes the constellation points. We saved the I-Q data obtained as the output of the Constellation Decoder block. We did not use any channel estimation techniques, in order not to filter out any channel effects useful for jamming detection. Regarding the jammer, we chained two blocks: (i) a signal source, which can be an analog $\sin$ wave (tone jammer) or a digital sequence of Gaussian-distributed values (Gaussian jammer); and (ii) the USRP Sink block, which sends the signal to the radio for actual transmission. To emulate the scenario described in Section III, we placed the entities at different distances and, to further emulate the movement, we changed the relative values of the jammer transmission power between 00 and $0.8$, i.e., between 0 and 7.94mW (9dBm), respectively. Values greater than $0.8$ cause a complete disruption of the BER of the signal (see Section V-C), making our solution unnecessary. In fact, when using a static setup, the reduction of the transmission power of the jammer makes the received jamming power level at the receiver weaker, allowing us to investigate the effect of a jammer located further away from the communication link.
For the second setup using the LimeSDR, we considered different values of sample rate at the receiver and at the jammer. In particular, consider the formula $t_{s}=K\times t_{S,R}$, where $t_{S,R}$ is the reference sample rate of 1M samples per second, $t_{s}$ is the actual sample rate used in the measurement, and $K$ an oversampling factor. We carried out experiments considering different values of the oversampling ratio both at the receiver and at the jammer, namely, the Receiver Oversampling Ratio (ROR) and Jamming Oversampling Ratio (JOR). Specifically, we varied both the ROR and the JOR in the range $[1,4]$. For all such experiments, we tested two jamming strategies: tone jamming, i.e., jamming with a sinusoid signal, and deceptive jamming, i.e., jamming with exactly the same bit-sequence BPSK-modulated signal delivered between the legitimate transmitter and the receiver.
Overall, the two measurement setups described above allowed us to investigate the effectiveness of BloodHound$+$ while varying a large range of configuration parameters including: (i) the RJP at the receiver, (ii) the distance of the jammer from the legitimate communication link, (iii) the oversampling ratios at the receiver and the jammer, and (iv) the types of jamming and the type of radios used for the experiments.
V-B Experimental Settings
For our experiments, we found the best configuration of the hyperparameters of the autoencoder of BloodHound$+$, and we compared its performance against the approach based on the binary CNN Resnet-18 used in [7] and the solution based on 3-D images proposed in [8].
Autoencoder deployment. We fine-tuned the hyperparameters of the autoencoder used in BloodHound$+$to find the best trade-off between classification accuracy and the general validity of the solution, i.e., to avoid overfitting. Specifically, we used the Matlab-provided implementation of autoencoders, version R2022b. As mentioned in Section IV, we used the Pure Linear (purelin) transfer function as the decoder transfer function. For all the remaining hyperparameters of the autoencoder, we used a formal hyperparameter optimization method conducted on a subset of our dataset. Specifically, we selected $1,500$ images from I-Q samples acquired in the scenario with the most data available, i.e., with the receiver positioned $10$ meters from the transmitter and the jammer emitting jamming with a RJP of $0.5$. We evaluated all combinations of the following hyperparameters: (i) hidden size (i.e., the size of the latent representation), with considered values being $8$, $16$, $32$ and $64$; (ii) sparsity regularization, with considered values being $1$, $0.5$, and 00; (iii) L2-regularization term, with considered values being $0.01$, $0.001$, and $0.0001$; and finally, (iv) encoder transfer function, with considered values being logsig and satlin. Finally, we set all remaining hyperparameters to their default values provided by Matlab, specified at [19]. Such a variety of configuration parameters led us to test a total number of $72$ configurations. For such tests, we used the methodology described below, inspired by the $k$-fold cross-validation technique used by the authors in [20].
  • We divide all jammed images in the selected dataset into $10$ disjoint subsets of equal size.
  • We divide all unjammed images in the selected dataset into evenly sized disjoint subsets of $10$.
  • For each considered combination of hyperparameters $\alpha$, for every $i\in\{1,\dots,10\}$, we do the following:
    • We train an autoencoder on all subsets of unjammed images, except for the $i$-th one. During training, we use the hyperparameters in $\alpha$.
    • We compute the mse values of the autoencoder on the images in the $i$-th subset of unjammed images.
    • We compute the mse values of the autoencoder on the images in the $i$-th subset of jammed images.
    • Using the mse values collected as part of the two steps above, we compute the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve corresponding to the autoencoder.
  • For each combination of hyperparameters $\alpha$, we take the average of the AUC-values found in the iterations of the last cited step.
  • Finally, we pick the combination of hyperparameters for which the last step mentioned above yielded the largest result.
Note that, like the authors in [20], we selected the AUC as the optimization metric, since it measures the quality of a classifier independently of the threshold selection process. Following such a hyperparameter selection process, we set the hidden size value to $16$, the sparsity regularization term to $0.5$, the L2-regularization term to $0.01$ and the encoder transfer function to the Logistic Sigmoid (logsig). We also selected the number of epochs by examining how many iterations the autoencoder takes to reliably converge in the worst case, i.e., when setting the size of the latent representation to 64. We empirically established this to occur after 250 epochs, and thus, we selected such a value to trade off between classification accuracy and training time. We use such a hyperparameter configuration to train the autoencoder of BloodHound$+$ for all the results discussed below on unseen data, so as to avoid overfitting.
Benchmark Approaches. For all experiments, we compare BloodHound$+$ with two benchmark solutions, i.e., the former version of BloodHound$+$ published in [7] and the approach proposed by the authors in [8]. We selected such solutions as they work successfully on PHY data, i.e., I-Q samples, significantly outperforming other solutions in challenging scenarios, such as the one considered in this manuscript. Both approaches use the residual CNN Resnet-18, pre-trained on the ImageNet dataset [21] and with the necessary modifications to the output layer necessary to fit the nature of the jamming detection problem. Specifically, the output layer of CNN has been modified to consider two possible classes as output, i.e., either No-Jamming or Jamming. Regarding the input layer, CNN works on images of size $224\times 224$ constructed over I-Q samples collected from the wireless channel, which is optimal for comparison to BloodHound$+$. To train CNN, we used the automated procedure trainNetwork provided by Matlab. We set the batch size to $32$, the number of epochs to $1$, and the solver to adam, similar to [7]. All the remaining hyperparameters have been set to their default values, available at [22].
The solution provided in [7] works on grayscale images, while the proposal in [8] uses color images, characterized by three layers rather than one. To set up such images, we follow the same procedure as described in [8]. As an example, Fig. 6 shows the comparison of a grayscale and a color image generated over the same set of I-Q samples collected from the wireless channel.
Figure 6: Sample grayscale and color unjammed (left) and jammed (right) images generated from the same set of I-Q samples, used for the experimental assessment. The jammed images have been obtained using samples collected at $10$ meters from the jammer, with RJP = $0.5$.
Measurements Characterization. As an introduction to the presentation of our results, with reference to the setup using the USRP X310 SDRs, in Fig. 7 we show the BER of the TX-RX communication link experienced with different values of the RJP, under different jamming signals.
Figure 7: Analysis of the BER of the TX-RX communication link using the SDRs USRP X310, with various levels of RJP and two jamming signals, i.e., tone jamming and Gaussian jamming. The receiver has been placed $10$ m away from the jammer.
We highlight that, when $RJP<0.7$, the impact of the jammer on the communication link in terms of BER is very limited, i.e., very few bits are corrupted. On the contrary, when $RJP\geq 0.7$, most of the bits are corrupted. Recall that in this manuscript we are specifically interested in early jamming detection in a mobile scenario, i.e., detecting the jammer before its impact on the communication link becomes significant and significantly affects the throughput of the communication link. In this context, we are particularly interested in enhancing jamming detection performances when $RJP<0.7$. For higher values of $RJP$, other techniques based on the analysis of BER can already detect jamming ex-post, i.e., once the communication link is significantly affected.
Finally, in this context, we note that the setup using the LimeSDR was configured with an absolute transmission gain at the jammer of $23$ dBm. This configuration allows us to have a BER of the legitimate communication link of approximately $0.001$, which allows us to study the impact of different configuration parameters of the scenario while matching the conditions of early jamming described above.
Performance Metrics. We compare the performance of the methodologies introduced above mainly with reference to their accuracy, obtained as $acc=\frac{TP+TN}{TP+FP+FN+TN}$, being $TP$ the amount of true positives (i.e., jammed images correctly classified), $TN$ the true negatives (i.e., unjammed images correctly classified), $FP$ the false positives (i.e., unjammed images wrongly classified as jammed ones) and $FN$ the false negatives (i.e., jammed images wrongly classified as unjammed ones). For some of the results shown below, we show the True Positive Ratio (TPR) and True Negative Ratio (TNR), computed as $TPR=\frac{TP}{TP+FP}$ and $TNR=\frac{TN}{TN+FN}$, respectively. We obtain our estimates of accuracy, TPR, and TNR using the cross-validation approach used by the authors of [20], using 10-fold cross-validation. For each result, we report the mean and the $95\%$ confidence intervals, computed using the tool tinv provided by Matlab.
V-C Early Jamming Detection Robustness
We first consider the impact of the received jamming power at the receiver on the capability of BloodHound$+$ to detect the presence of the jammer. We consider the setup using the Ettus X310 SDR, and specifically, the measurements where the receiver is located $10$ meters away from the jammer. Here, we generate images using $n=10^{5}$ I-Q samples and evaluate the performance of BloodHound$+$, the proposal in [8] and the solution in [7] to detect jamming, in terms of overall classification accuracy. We report the results of our investigation in Fig. 8.
Figure 8: Classification accuracy of BloodHound$+$, the proposal in [8] and the solution in [7] with various levels of RJP.
First, we note that the approaches in [8] and [7], based on CNNs, reliably identify jamming only when $RJP\geq 0.4$. With lower values of the RJP, their performances do not follow a unique trend, and the results also exhibit high variance (see the red bars indicating the confidence intervals of the measurements). On the contrary, BloodHound$+$reports remarkable performances for every value tested of $RJP$, with a minimum accuracy of $0.997$. We believe that such a result is due to the rationale of autoencoders used in BloodHound$+$, which work only on unjammed images. Such a configuration allows autoencoders to build a profile of the regular behavior of the wireless channel, in order to reliably identify smaller differences (anomalies).
We also investigated the impact of the distance of the receiver from the jammer on the performance of the cited solutions. To this aim, using the same setup as used for the previous tests, we set $RJP=0.5$ and move the receiver away from the jammer, up to a distance of $21$ meters. We stop at such a distance due to the physical limitations of the involved hardware, i.e., at distances higher than $21$ meters the BER of the legitimate communication link decreased significantly, preventing us from performing the test reliably. We report the results of our experiments in Fig. 9.
Figure 9: Classification accuracy of BloodHound$+$, the proposal in [8] and the solution in [7] when positioning the receiver at various distances from the jammer.
Here, we notice that all the tested solutions report remarkable performance, with very limited differences. All three approaches reliably detect jamming at various distances, with an average accuracy of well above $0.99$. Overall, the distance between the receiver and the transmitters does not affect the classification accuracy, which remains very high even when the receiver is $21$ meters away.
Another important parameter of BloodHound$+$ is the number of samples used to generate images, namely $n$. The higher the value of such a parameter, the higher the number of samples to use for image generation, thus the longer the receiver has to acquire samples from the wireless channel, and the higher the processing overhead of the solution. For this analysis, in the same setup as in the previous experiments, we considered the data acquired with the receiver located $10$ m away from the jammer and $RJP=0.5$, and tested the performance of the three approaches while increasing the number of samples used in the image generation phase, from $10,000$ to $1,000,000$. Fig. 10 reports the results of our analysis.
Figure 10: Classification accuracy of BloodHound$+$, the proposal in [8] and the solution in [7] considering an increasing number of samples $N$ in the Image Generation process.
Note that the solutions in [8] and [7] report a higher classification accuracy than BloodHound$+$ for a low number of samples. For example, when the $n=10,000$, BloodHound$+$ reports an accuracy of $0.498$ while such values amount to $0.99$ and $0.95$ for the solution in [8] and [7], respectively. When the available number of samples increases, the accuracy of the benchmark approaches is still high, but the variance becomes larger (see the red bars). Conversely, when $n\geq 50,000$, not only the accuracy of BloodHound$+$ is very high (always higher than $0.99$), but the variance is also very limited (less than $0.001$ for all tests), which indicates greater robustness and reliability. Such results further motivate the deployment of BloodHound$+$ and highlight its superiority compared to the benchmark solutions.
To provide further insight into the performance of BloodHound$+$, we investigated the impact of additional configuration parameters. In particular, in the same setup as the last cited test, considering $n=100,000$, we evaluated the impact of the size of the training set. Figure 11 summarizes the results of our investigation, distinguishing the achieved TPR and TNR of our proposed solution.
Figure 11: TPR and TNR of BloodHound$+$ when varying the training set size, i.e., the number of unjammed images used for training the autoencoder.
We notice that to perform reliably, BloodHound$+$requires a minimum training set of only $9$ images, reporting $TPR$ and $TNR$ values of $0.962$ and $0.99$, respectively. Performance stays almost constant when increasing the training set size. However, increasing the training set could be particularly relevant in very noisy scenarios, characterized by a wider range of expected wireless channel fluctuations. We recall that we need to train BloodHound$+$ only once, before deployment, and such results do not affect the deployability of our solution at runtime.
We also investigated further any bias of our results with respect to the specific hardware used for the experiments. Taking into account the same scenario as in previous experiments, i.e., the receiver located $10$ meters away from the jammer, we evaluated the TPR and TNR of BloodHound$+$ when changing the hardware used for jamming among the five available radios. We highlight that this methodology prevents the autoencoder from fingerprinting both the transmitter and the receiver, these being the same for all measurement classes. Moreover, we considered different hardware for the jammer during our measurements, i.e., we mutually excluded the ones adopted for training from the ones adopted for testing. This eventually guarantees that the autoencoder learns the characteristics of the legitimate signal only, while being independent of the transmitter, the receiver, and the jammer hardware (jammed signals are not used for training). Specifically, we consider all unjammed images obtained when placing the receiver at a distance of 10 meters from the transmitter, using radio $x$ as the jammer. Next, we considered all jammed images generated with the receiver located at a distance of 10 meters and using radio $y$ as a jammer. We separated the unjammed and jammed images into 10 folds, and we trained BloodHound$+$ on 9 of the folds containing unjammed images, holding the $i$-th one out to estimate the TNR. Next, we estimate the TNR on the $i$-th fold of unjammed images and estimate the TPR on the $i$-th fold of jammed images. Fig. 12 reports the result of our analysis. The tick labels on the x-axis are of the form $(x,y)$, as described above.
Figure 12: TPR and TNR of BloodHound$+$ when varying the hardware used for jamming and the radio used for training.
Note that the TPR and TNR remain almost unchanged while varying the hardware used for jamming and the radio considered for training. Therefore, we can safely assume that BloodHound$+$ is not biased by the specific radios used in the experiments, but it can extract the features of the wireless channel useful for detecting jamming, being independent of specific hardware.
V-D Impact of Different Hardware and Sampling Rate
The results shown in the previous subsection have all been obtained using a single hardware brand (Ettus Research USRP X310) and a single configuration of the sample rate, i.e., $1$ Msa/s. To further extend our experimental assessment, we first ran further analyzes on the data collected through the second setup described in Section V-A, adopting the hardware LimeSDR Mini. An important consideration about the scenario described in Section III is that the jammer might not know in advance what the adopted sampling rate is. Thus, to disrupt ongoing communications as much as possible, in real-life scenarios, the jammer might emit jamming using the maximum achievable sampling ratio, likely higher than the one adopted by the legitimate communication link. At the same time, the receiver might oversample the signal, obtaining more useful information for jamming detection. This additional information might be discarded for communication purposes, but it might still be useful for jamming detection. Therefore, in our tests, we trained BloodHound$+$ on unjammed images of the legitimate communication link obtained under $RJP=0.5$ and a distance of 3 meters, and we tested using unjammed images (disjoint from the ones used for training) and jammed images obtained from I-Q samples generated with various levels of JOR. For this test, we let the jammer emit random Gaussian noise. We also compared the performance of BloodHound$+$ with the ones of the solution in [7]. Fig. 13 summarizes the results of our analysis, in terms of TPR (for comparison purposes).
Figure 13: True Positive Ratio (TPR) of BloodHound$+$ and the proposal in [7], with various ROR and JOR values, when the jammer injects Gaussian noise (Gaussian random jamming).
Note that the TNR of BloodHound$+$ was also excellent, at an average of 0.987 over all values of JOR.
We can distinguish two effects from the reported results. Taking into account the configuration with $ROR=1$ and $JOR=1$, we notice that the solution in [7] already reports very low TPR ($0.03$). We highlight that this result is not related to the usage of different oversampling rates but to the new hardware used for the experiments. Indeed, the LimeSDR is a cheaper hardware, which introduces additional inaccuracies and imperfections in the I-Q samples received. Such inaccuracies affect the shape of the I-Q samples, which is now more spread around the expected symbol than before, leading to jammed images different from the ones in the training set of the solution in [7]. These variations are evidently not caught by the approach in [7], which fails to perform reliable jamming detection. On the contrary, in such a configuration, the performance of BloodHound$+$ does not change compared to the results shown in Section V-C, demonstrating once again the enhanced robustness offered by the autoencoders used in BloodHound$+$ compared to the CNNs used in the competing solution.
With higher values of ROR and JOR, the performance of the approach in [7] remains well below $0.2$, confirming the unsuitability of such a solution for the detection of jammers in the wild. Instead, BloodHound$+$ can mitigate and overcome the impact on the wireless channel of different oversampling ratios, being able to detect jamming also when the JOR is very high.
V-E Deceptive Jamming
In previous experiments, we mainly considered two jamming models, that is, tone jamming (using a sinusoid signal) and random jamming (using AWGN). When the adversary does not know the modulation used by the legitimate communication link, the jamming models mentioned and investigated above are the most reasonable options to disrupt the channel as much as possible. However, more powerful attackers might know in advance or become aware at runtime (e.g., by eavesdropping) of the modulation used by the legitimate communication link. Based on such knowledge, they might use the signal as part of a jamming attack optimized to disrupt the communication link as much as possible. Recall that in our experiments the legitimate communication link adopts the BPSK modulation scheme. In such a scenario, the usage of a jamming signal modulated also as a BPSK allows to boost the effectiveness of the jamming activity, as acknowledged by several scientific contributions both analytically [23] and experimentally [24].
To this end, using the same setup as the previous experiments in Section V-D, we evaluated the ability of both BloodHound$+$and the solution in [7] to detect deceptive jamming, i.e., a jammer injecting the same signal used as part of the legitimate communication link. Fig. 14 summarizes the results of our analysis in terms of TPR (for comparison purposes). Note that, also in this case, the TNR of BloodHound$+$ is 0.987, on average over all JOR, since we trained the autoencoder in BloodHound$+$ with the same data as in Fig. 13.
Figure 14: True Positive Ratio (TPR) of BloodHound$+$ and the proposal in [7], with various ROR and JOR values, when the jammer injects the same signal of the legitimate communication link (deceptive jamming).
Even when the adversary uses deceptive jamming, BloodHound$+$ significantly outperforms the benchmark solution in all tested configurations, showing perfect TPR and being also robust with respect to the adoption of a high JOR by the adversary.
Overall, the results reported above demonstrate the superiority of BloodHound$+$ compared to benchmark approaches, as well as the robustness of our solution to a wide range of configuration parameters and scenarios, making it the preferred solution for early jamming detection.
VI Related Work
Several scientific papers recently considered drones for indoor applications, focusing on aspects such as localization [25], navigation [26], [27], and visualization [28]. However, none of them investigates jamming attacks and anti-jamming approaches for indoor scenarios, thus mainly referring to the literature on generic (outdoor) jamming detection.
In the scientific community, jamming detection is usually achieved by applying various types of analysis on one or more metrics extracted from the primary communication link. Regarding the metrics, several parameters have been analyzed, such as the RSS of the signals [29], the PDR [30] as in [31], the profile of packet retransmissions, as in [32], or modulation-specific metrics, such as for Orthogonal Frequency Division Multiplexing (OFDM) in [33].
At the same time, due to the increasing popularity of Artificial Intelligence (AI), Machine Learning (ML) and DL approaches have been recently used extensively for detecting ongoing jamming. Such tools include CNNs such as in [34] and [35], genetic algorithm-based cumulative sum methods such as in [36], Bayesian networks such as in [37] and, finally, autoencoders such as in [38]. All such approaches utilize as the main source of information the PHY layer of the communication stack, due to its direct relationship with the wireless channel, where jamming occurs. However, although some of the cited contributions analyzed the performance of the proposed jamming detection technique with low SNR, none of them considered the BER of the communication link. As a result, the proposed approaches mostly confirm that the root cause of the drop in BER is jamming. However, they cannot detect such attacks even when the jamming effect is so low as not to affect the BER of the communication link. As explained above, such consideration is particularly relevant in mobile scenarios for remotely controlled equipment, not to lose control of the mobile entity completely before detecting jamming. In this context, the only contribution in the literature achieving such a property is our previous proposal in [7]. As shown in Section V the methodology shown in this paper significantly outperforms both the one proposed in [7] and the evolutions of such a methodology as the one proposed in [8].
VII Conclusion
In this paper, we have presented BloodHound$+$, an approach that allows drones and possibly other mobile devices to detect jamming at the physical layer (PHY) of the communication stack. Our solution works on raw I-Q samples extracted from the communication link, converts them into grayscale images, and uses sparse autoencoders to detect discrepancies with the expected profile of the channel. Therefore, BloodHound$+$ can detect jamming well before its effect could cause a significant decrease in the quality of the main communication link. In turn, our solution allows drones to efficiently avoid the jammed area and maintain full control and safety. To test the effectiveness of our solution, we conducted an extensive measurement campaign, acquiring real-world data with different hardware, jamming strategies, and scenario configurations. We also tested the performance of BloodHound$+$ depending on various parameters, such as SNR of the communication link, the distance from the jammer and the transmitter, the size of the training set, the number of samples acquired from the channel, the jammer oversampling ratio, and the jamming strategies. Our experimental assessment demonstrates, through a variety of results, the superiority of our solution compared to the current state-of-the-art across all the analyzed configuration parameters.
In general, we believe our solution contributes to taking a step further toward the safe and secure integration of drones into daily life. As part of our future work, we plan to further investigate the effectiveness of BloodHound$+$ when applied for outdoor applications.
Acknowledgements
This work has been partially supported by the INTERSECT project, Grant No. NWA.1162.18.301, funded by the Netherlands Organization for Scientific Research (NWO). Any opinions, findings, conclusions, or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of NWO. Moreover, this research was sponsored in part by the NATO Science for Peace and Security Programme under grant SPS G5797.